Dear Richard dear all,

please find below a short summary of the activities of 13th and 14th on the mountain related to the AO system.
Italian people working at LBTO were:

March 13th

Arcetri: P. Stefanini, A. Riccardi, R. Briguglio, L. Busoni, M. Xompero
ADS: P. Lazzarini
LBTO: G. Brusa

Total # of person: 7

W unit

1) Remote functionality test of the unit was repeated successfully. A ROn test with the dark dome has been done with the following results:
2.5 Mpix/s 8.60, 860 Kpix/s 5.82, 400 Kpix/s 4.07. These results are better than those obtained in the lab in Arcetri during acceptance test.
The improvement is of about 2 electrons for the 2.5 Mpix/s and 0.5 electrons for the 860 Kpix/s.


AO control SW

1) Debug of control SW for non corrected modes.
2) Debug of control SW for reference fibers located on the secondary mirror.

LBT672a

1) The ASM cover has been removed and the retroreflector holder has been installed on the telescope. The procedure for this installation has been adapted for the telescope configuration in which the secondary is deployed. The task showed some criticality that will require some modifications of the HW.
2) A functional check of the unit after the retroreflector installation has been done showing the following
- 4 contiguous actuators of the external ring have a malfunctioning behavior compatible with the grounding of the capacitive sensor contact.
- The setting of the mirror is no longer possible below 10 degree of elevation because of overcorrect threshold.
3) Night time test were run for the first time with the dome shutter of the DX side opened and the shell exposed to the wind without the cover. No proble were detected with a wind gust up to 8 m/s.
4) The interferometer laser was found to have unstable light output. The laser unit had a broken connector. A preliminary alignment with the unit with M3 and M2 was attempted.


March 14th

Total # of person: 7

Arcetri: E. Pinna, A. Riccardi, R. Briguglio, M. Xompero, A. Puglisi, S. Esposito
LBTO: G. Brusa


LBT672a

1) mirror flattening. The task has been interrupted to investigate the interferometer failure. the interferometer has been moved back to the clean lab. After contacting 4D company it was decide to replace the laser unit. The new unit is expected to be on the mountain by tomorrow midday. After the work of today and the present status of the interferometer the possibility to have the flattening task completed before the AGw commissioning is quite low.

FLAO activity

1) Because of the problem with the interferometer it was decided to start the alignment of the complete AO system. This in order to possibly flat the ASM using the WFS. The M3 unit was moved in the position found by A. Rakich in its alignment campaign. The M3 unit has been operated by P. Lazzarini before leaving the observatory. No other movement of M3 was possible during the day. After some attempts using the polychromatic source of the W unit we installed on the AGW our laser to project a more intense light beam. The light beam has been seen focused by the secondary mirror on the top of the retroreflecting optics. This was done by using a binocular from the 9th floor. The spot has been located and the alignment process will be restarted tomorrow when M3 operation will be possible.



Next update tomorrow

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback